OpenStudy (anonymous):

evaluate the limit lim x->5

OpenStudy (solomonzelman):

what is the limit ?

OpenStudy (anonymous):

\[\lim_{x \rightarrow 5} \frac{ |x^2-4x-5| }{ x^2-25 }\]

OpenStudy (anonymous):

can u show me step by step?

OpenStudy (dan815):

rewrite the top abs value

OpenStudy (solomonzelman):

factor \(\normalsize\color{midnightblue}{ x^2-25=(x-5)(x+5) }\) and factor out the top \(\normalsize\color{midnightblue}{ │x^2-4x-5│=│(x-5)(x+1)│ }\) and then despite absolute value divide top and bottom by \(\normalsize\color{midnightblue}{ (x-5) }\)

OpenStudy (solomonzelman):

then plug in 5 instead of x.

OpenStudy (anonymous):

but it is 1/2 why is answer dne?

OpenStudy (solomonzelman):

the answer is not 1/2

OpenStudy (anonymous):

-1/2?

OpenStudy (solomonzelman):

What do you get after factoring top and bottom and dividing top and bottom by (x-5 ) (without removing the absolute value from the top ) ?

OpenStudy (anonymous):

1/2?

OpenStudy (solomonzelman):

it really seems like you are trying to go into calc without knowing even simple algebra.

OpenStudy (solomonzelman):

I am out, jim thompson will help

jimthompson5910 (jim_thompson5910):

make a table and you'll find that as x approaches 5 from the left, the limit is different than if you were to approach 5 from the right So because the left hand limit and the right hand limit are NOT equal, this means the limit as x --> 5 does NOT exist

OpenStudy (anonymous):

i think i understand. i think i have to find the values though i cant just say dne?

OpenStudy (anonymous):

what if I have \[\lim_{x \rightarrow \infty} \frac{ \cos^5(2x-1) }{ x^3 }\]??

jimthompson5910 (jim_thompson5910):

Here's the graph and table I'm getting. I used geogebra to get all this.

jimthompson5910 (jim_thompson5910):

Notice the jump discontinuity at x = 5

OpenStudy (anonymous):

oh i see that

jimthompson5910 (jim_thompson5910):

and notice how the table reflects in this large jump/gap you should be approaching a single value as you approach 5, but that's not the case

jimthompson5910 (jim_thompson5910):

what's the range of cosine?

OpenStudy (anonymous):

-1 to 1

jimthompson5910 (jim_thompson5910):

so cos^5(x) will also have a range from -1 to 1 because (-1)^5 = -1, (1)^5 = 1

jimthompson5910 (jim_thompson5910):

basically, no matter what the argument does, the numerator is going to bounce between -1 and 1

jimthompson5910 (jim_thompson5910):

as x ---> infinity, the denominator grows bigger but the numerator is fixed to between -1 and 1

jimthompson5910 (jim_thompson5910):

so the whole fraction gets so very very small that it approaches 0

OpenStudy (anonymous):

oh right. are you a teacher u teach better than mine!!

jimthompson5910 (jim_thompson5910):

I'm studying to become one I'm glad it's clicking a bit better now

OpenStudy (anonymous):

yes. i have a quiz tomorrow. can i ask just one more?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

is it same thing for this one? \[\lim_{x \rightarrow 0}x^6\cos^5(\frac{ 1 }{ 5x}- \pi)\]

jimthompson5910 (jim_thompson5910):

yeah no matter what, cos is going to bounce around from -1 to 1 as x --> 0, x^6 ---> 0 since x^6 = 0^6 = 0

jimthompson5910 (jim_thompson5910):

so we really have 0*cos(x) where x is some number, we really don't need to worry about what its value is

jimthompson5910 (jim_thompson5910):

and that whole thing is 0

OpenStudy (anonymous):

if they put sine then answer is also same?

OpenStudy (anonymous):

thank you i understand tshis a lot better. wow.

jimthompson5910 (jim_thompson5910):

yep because they both have the same range

jimthompson5910 (jim_thompson5910):

it really applies to any function that has a restricted range